OpenStudy (anonymous):

from the top of a tower of height 20m a boy drops one stone ,one second later he throws another stone vertically downwards so that both hit the ground together the initial velocity of the second stone is (g=10m/s^2)

OpenStudy (ajprincess):

try this and tell me if you find any difficulties

OpenStudy (anonymous):

i tried this but i got different answer so i posted this question...... all the question which i am posting is already tried by me ......

OpenStudy (ajprincess):

ohh k.

OpenStudy (ajprincess):

let me guide u

OpenStudy (anonymous):

thats a good idea....

OpenStudy (anonymous):

i think first we have to find time???

OpenStudy (ajprincess):

nw first give me the answer?

OpenStudy (anonymous):

15m/s

OpenStudy (ajprincess):

As usual use s=ut+1/2at^2

OpenStudy (anonymous):

we have to find t is it....

OpenStudy (ajprincess):

ya u r right.

OpenStudy (anonymous):

t=2s

OpenStudy (ajprincess):

this one u worked out nw or the given answer?

OpenStudy (anonymous):

next what to do t1=t2

OpenStudy (ajprincess):

the times wont be equal.

OpenStudy (ajprincess):

t1=t2-1

OpenStudy (anonymous):

yes u r right

OpenStudy (ajprincess):

for jst dropping the initial velocty is 0.

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

so u cn find t1 using that adn t2 too.

OpenStudy (anonymous):

t2=3s

OpenStudy (ajprincess):

jst a moment

OpenStudy (ajprincess):

sorry t1=t2+1. so t2=1s

OpenStudy (anonymous):

oh now got it..........thanzzz

OpenStudy (ajprincess):

yww